A lab was interested in studying the average weight of mice after being given a specific diet. The following table includes the weight of a small sample of the mice after the diet period. Mouse 1 10g, Mouse 2 8g, Mouse 3 11g, Mouse 4 10g, Mouse 5 9g, What is the 95% confidence interval for the mean weight of all lab mice?

(a) 9.6 ± 2.7764(1.140/ √ 5)

(b) 9.6 ± 1.96(1.140/ √ 5)

(c) 9.6 ± 2.0150(1.140/ √ 5)

(d) 9.6 ± 2.1318(1.3/ √ 5)

(e) 9.6 ± 2.0150(1.3/ √ 5)
Answer #1

= (10 + 8 + 11 + 10 + 9)/5 = 9.6

s = sqrt(((10 - 9.6)^2 + (8 - 9.6)^2 + (11 - 9.6)^2 + (10 - 9.6)^2 + (9 - 9.6)^2)/4) = 1.14

At 95% confidence level, the critical value is t0.025, 4 = 2.7764

The 95% confidence interval is

Option - a is correct.